I need to solve two different logarithmic equations.

1) 2log10x=10 (the first 10 is the base ten)
2) 9 log 10 x-4 = 11 (the ten is also the base in this equation, and the x - 4 is the exponent)

Answer by dfrazzetto(283)

If the base is ten, just write log, it's understood:
2logx = 10
log(x) = 5
10^5 = x
x = 100,000


9*log(x-4) = 11
log(x-4) = 11/9
x-4 = 10^(11/9)
x = 10^(11/9) + 4
x = 16.68 + 4 = 20.68